# Automated Segmentation of Coronal Brain Tissue Slabs for 3D Neuropathology

### v1

@@ Abstract
Advances in image registration and machine learning have recently enabled volumetric analysis of postmortem brain tissue from conventional photographs of coronal slabs, which are routinely collected in brain banks and neuropathology laboratories worldwide. One caveat of this methodology is the requirement of segmentation of the tissue from photographs, which currently requires costly manual intervention. In this article, we present a deep learning model to automate this process. The automatic segmentation tool relies on a U-Net architecture that was trained with a combination of (i)1,414 manually segmented images of both fixed and fresh tissue, from specimens with varying diagnoses, photographed at two different sites; and (ii) 2,000 synthetic images with randomized contrast and corresponding masks generated from MRI scans for improved generalizability to unseen photographic setups. Automated model predictions on a subset of photographs not seen in training were analyzed to estimate performance compared to manual labels – including both inter- and intra-rater variability. Our model achieved a median Dice score over 0.98, mean surface distance under 0.4 mm, and 95% Hausdorff distance under 1.60 mm, which approaches inter-/intra-rater levels. Our tool is publicly available at surfer.nmr.mgh.harvard.edu/fswiki/PhotoTools.
@@ Introduction
Neuropathologists now have an growing array of tools for tissue analysis, ranging from conventional histological stains [11] to advanced molecular techniques, including RNA sequencing and multi-omics approaches such as spatial transcriptomics applied to frozen tissue [12]. There is growing interest in integrating these molecular and histopathological data with neuroimaging modalities obtained antemortem (e.g., magnetic resonance imaging, computed tomography, or positron emission tomography) to identify imaging biomarkers that could ultimately enable diagnosis and monitoring in vivo [13, 14].
Magnetic resonance imaging (MRI) is widely regarded as the gold standard for morphometric analysis of brain specimens due to its high soft tissue contrast, excellent spatial resolution, and non-invasive nature. However, antemortem MRI is often unavailable or acquired long before death, during which time substantial anatomical changes may occur, limiting its utility for direct comparison with postmortem histopathology. Cadaveric MRI presents a viable alternative, offering the ability to scan intact brains shortly after death. However, it requires rapid acquisition prior to tissue degradation that could compromise downstream analyses [15]. Such capability is typically restricted to specialized facilities equipped with dedicated research scanners and appropriate infrastructure. An additional option is ex vivo MRI, which provides greater flexibility in scheduling but necessitates advanced expertise and customized protocols that may not be readily available [16]. Despite these advances, most brain banks and neuropathology laboratories worldwide lack access to MRI scanning resources altogether. While positron emission tomography (PET) enables functional imaging capabilities not possible with MRI, it is restricted to in vivo use and remains limited in availability.
@@
A major challenge in scaling 3D analysis of dissection photography is the need for brain tissue segmentation, which currently requires labor-intensive manual annotation. This is primarily due to the presence of the cortical surface in the images (Figure 1), which is not part of the slab face and should be excluded from downstream analyses [13]. Segmenting these non-relevant regions is particularly time-consuming, as conventional thresholding techniques (most often effective against high-contrast backgrounds) fail to accurately distinguish the cortical rim from relevant tissue. As a result, manual delineation is often necessary. Automating this segmentation step would significantly reduce the manual workload and enhance the scalability and reproducibility of the pipeline, making high-throughput analysis of dissection photographs more feasible.
@@
We trained our model using 1,414 real photographs from three sources, comprising fresh and fixed tissue. Our experiments show that the model achieves segmentation performance on par with manual annotations. Furthermore, even when manual corrections are necessary, the automated predictions substantially reduce the time required by expert labelers, improving overall labeling efficiency. Importantly, the segmentation tool is publicly available and has been integrated into the FreeSurfer software suite [22], facilitating broad adoption within the neuroimaging and neuropathology communities.
@@
Deep learning has rapidly advanced computational neuropathology. Convolutional neural networks (CNNs [30, 31]) and more recently transformers [32] enabled accurate and scalable solutions to tasks like tissue segmentation, cell detection, or cell-type classification. In segmentation, architectures like the ubiquitous U-Net [20] enable automatic delineation of tissue compartments or nuclei with high accuracy. Detection models, including region-based CNNs, focus on identifying individual cells or pathological structures, while classification models typically operate on image patches or single-cell representations to assign cell types or disease labels [36]. More recently, foundation models have introduced new levels of generalizability. Promptable segmentation models such as SAM [37] and MedSAM [38] can adapt to a wide range of biomedical images with minimal task-specific training. In digital pathology, large-scale pretrained models have recently demonstrated strong generalizability and performance across tasks [39].
@@ Methods
Same segmentations are shown in Figure 3 and in the supplement Figure S1 to S6, whereas quantitative results are shown in Figure 4. On in-distribution data, the model achieved excellent performance, with a median Dice score of 0.985, a mean surface distance of 0.34 mm, and a HD95 of 1.27 mm. These values closely approach the intra-rater variability of the more consistent human annotator (noting that this comparison involves different image samples). Using a very conservative threshold of 2 mm for the mean surface distance to define outliers, the in-distribution outlier rate was just 2%.
@@ Discussion
This work introduces a fully automated segmentation tool for brain dissection photographs, offering a practical and scalable alternative to resource-intensive annotation workflows. By leveraging the nnU-Net framework and training on a combination of real and synthetically generated data, our method achieves accuracy comparable to that of expert human annotators. It maintains strong performance not only on in-distribution data, but also generalizes effectively to out-of-distribution images from a previously unseen site.
@@
To this end, future work may explore additional strategies to enhance robustness across domains. For example, synthetic data generation with domain randomization could further expose the model to a broader range of appearances and imaging conditions. Incorporating style transfer or domain adaptation techniques may also reduce sensitivity to site-specific artifacts. Another potential direction direction is to use lightweight image-level quality control (QC) mechanisms, such as binary pass/fail ratings, to complement traditional spatial metrics like Dice and surface distance. Such QC frameworks may provide more intuitive assessments of segmentation reliability in large-scale, real-world workflows.
Ultimately, the practical impact of this tool lies in its potential to dramatically reduce the time burden of manual annotation. As we estimated, segmenting a typical dissection photograph takes approximately 10 minutes by hand. Extrapolated to our training dataset alone, this represents over 200 hours of expert time. Our tool can reduce this workload to near-zero for most images and still substantially accelerate expert review when manual refinement is needed. This represents a significant step toward scalable, high-throughput analysis of neuropathology specimens, unlocking the potential of vast image archives that would otherwise be prohibitively time-consuming to label manually.
@@ Acknowledgments
This research was primarily funded by the National Institute of Aging (1R01AG070988). Additional support was provided by the National Institute of Mental Health (1RF1MH123195, 1UM1MH130981), the National Institute of Aging (1RF1AG080371), the National Institute of Biomedical Imaging and Bioengineering (1R01EB031114), and the National Institute of Neurological Disorders and Stroke (1R21NS138995). The UW BioRepository and Integrated Neuropathology (BRaIN) laboratory is supported by the National Institutes of Health (NIH) through the UW Alzheimer’s Disease Research Center (P30 AG066509), the Kaiser Permanente Washington Adult Changes in Thought (ACT) study (U19 AG066567), the BRAIN Initiative Cell Atlas Network (UM1MH134812), the Seattle Alzheimer’s Disease Brain Cell Atlas (U19AG060909), multiple collaborative projects supporting imaging, digital pathology, neuropathology, and related research tools (U24AG072458; U24NS133949; U24NS133945; U24NS135651; U01NS137500; and U01NS137484), the US Department of Defense (DoD W81XWH-21-S-TBIPH2), the Allen Institute for Brain Science, and the Nancy and Buster Alvord Endowment (to C.D.K.). The authors would like to thank the research participants and their families without whom this work would be impossible.
@@
Dr Hyman owns stock in Novartis; he serves on the SAB of Dewpoint and has an option for stock. He serves on a scientific advisory board or is a consultant for AbbVie, Alexion, Ambagon, Aprinoia Therapeutics, Arbor Bio, Arvinas, Avrobio, AstraZenica, Biogen, Bioinsights, BMS, Cell Signaling, Cure Alz Fund, CurieBio, Dewpoint, Etiome, Latus, Merck, Novartis, Paragon, Pfizer, Sanofi, Sofinnova, SV Health, Takeda, TD Cowen, Vigil, Violet, Voyager, WaveBreak. His laboratory is supported by research grants from the National Institutes of Health, Cure Alzheimer’s Fund, Tau Consortium, and the JPB Foundation – and sponsored research agreement from Abbvie and Sanofi. He has a collaborative project with Biogen and Neurimmune.

### v2

@@ Abstract
Advances in image registration and machine learning have recently enabled volumetric analysis of postmortem brain tissue from conventional photographs of coronal slabs, which are routinely collected in brain banks and neuropathology laboratories around the world. One caveat of this methodology is the requirement of segmentation of the tissue from the background and out-of-slice tissue in photographs, which currently requires laborious manual intervention. In this article, we present a deep learning model to automate this process. The automatic segmentation tool relies on a U-Net architecture that was trained with a combination of 1,414 manually segmented images of both fixed and fresh tissue, from specimens with varying diagnoses, photographed at two different sites. Automated model predictions on a subset of photographs not seen in training were analyzed to estimate performance compared to manual labels, including both inter- and intra-rater variability. Our model achieved a median Dice score over 0.98, mean surface distance under 0.4 mm, and 95% Hausdorff distance under 1.60 mm, which approaches inter-/intra-rater levels. Our tool is publicly available at surfer.nmr.mgh.harvard.edu/fswiki/PhotoTools.
@@ Introduction
Neuropathologists now have an increasing array of tools for tissue analysis, ranging from conventional histological stains [11] to advanced molecular techniques, including RNA sequencing and multi-omics approaches such as spatial transcriptomics applied to frozen tissue [12]. There is growing interest in integrating these molecular and histopathological data with neuroimaging modalities obtained antemortem (e.g., magnetic resonance imaging, computed tomography, or positron emission tomography) to identify imaging biomarkers that could ultimately enable diagnosis and monitoring in vivo [13, 14].
Magnetic resonance imaging (MRI) is widely regarded as the gold standard for morphometric analysis of brain specimens due to its high soft tissue contrast, excellent spatial resolution, and non-invasive nature. However, antemortem MRI is often unavailable or acquired long before death, during which time substantial anatomical changes may occur, limiting its utility for direct comparison with postmortem histopathology. Cadaveric MRI presents a viable alternative, offering the ability to scan intact brains shortly after death. However, unless the brain has been fixed, MRI requires rapid acquisition prior to tissue degradation that could compromise downstream analyses [15]. Such a capability is typically restricted to specialized facilities equipped with dedicated research scanners and appropriate infrastructure. An additional option is ex vivo MRI, which provides greater flexibility in scheduling but necessitates specialized expertise and customized protocols that may not be readily available [16]. Despite these advances, most brain banks and neuropathology laboratories worldwide lack access to MRI scanning resources altogether. While positron emission tomography (PET) enables functional imaging capabilities not possible with MRI, it is restricted to in vivo use and remains limited in availability.
@@
A major challenge in scaling 3D analysis of dissection photography is the need for segmenting brain tissue from the background and out-of-slice tissue, which currently requires labor-intensive manual annotation. This is primarily due to the presence of the cortical surface in the images (Figure 1), which is not part of the slab face and should be excluded from downstream analyses [13]. Segmenting these non-relevant regions is particularly time-consuming, as conventional thresholding techniques (most often effective against high-contrast backgrounds) fail to accurately distinguish the cortical rim from relevant tissue. As a result, manual delineation is often necessary. Automating this segmentation step would significantly reduce the manual workload and enhance the scalability and reproducibility of the pipeline, making high-throughput analysis of dissection photographs more feasible.
@@
We trained our model using 1,414 real photographs from three sources, comprising fresh and fixed tissue. Our experiments show that the model achieves segmentation performance on par with manual annotations. Furthermore, even when manual corrections are necessary, automated predictions substantially reduce the time required by expert labelers, improving overall labeling efficiency. Importantly, the segmentation tool is publicly available and has been integrated into the FreeSurfer software suite [22], facilitating broad adoption within the neuroimaging and neuropathology communities.
@@
Deep learning has rapidly advanced computational neuropathology. Convolutional neural networks (CNNs [30, 31]) and more recently transformers [32] enabled accurate and scalable solutions to tasks such as tissue segmentation, cell detection, or cell-type classification. In segmentation, architectures like the ubiquitous U-Net [20] enable automatic delineation of tissue compartments or nuclei with high accuracy. Detection models, including region-based CNNs, focus on identifying individual cells or pathological structures, while classification models typically operate on image patches or single-cell representations to assign cell types or disease labels [36]. More recently, foundation models have introduced new levels of generalizability. Promptable segmentation models such as SAM [37], MedSAM [38], and ScribblePrompt [39], can adapt to a wide range of biomedical images with minimal task-specific training. In digital pathology, large-scale pretrained models have recently demonstrated strong generalizability and performance across tasks [40].
@@ Methods
Ethics Statement
Human tissue used in this study was obtained in accordance with institutional and national ethical guidelines. Procedures for tissue collection and processing at Massachusetts General Hospital were reviewed and approved by the Partners Human Research Committee institutional review board under protocol 1999P009556. All postmortem specimens were obtained with informed consent from next of kin. At the University of Washington, tissue processing and collection involves the receipt and analysis of de-identified information and specimens from deceased individuals which were acquired from a biorepository. Consent for the use of information and specimens from the deceased individuals was obtained either from the individual while they are alive, or from the individual’s legally authorized representative. Procurement and banking practices of the biorepository are informed by the US Revised Uniform Anatomical Gift ACT 2006 (Last Revised or Amended in 2009) and Washington Statute Chapter 68.64 RCW. We work closely with the UW School of Medicine Compliance Office on our consent forms and HIPAA compliance. All materials are collected under informed consent.
@@
Some segmentations are shown in Figure 3 and in the supplement Figure S1 to S6, whereas quantitative results are shown in Figure 4. On in-distribution data, the model achieved excellent performance, with a median Dice score of 0.985, a mean surface distance of 0.34 mm, and a HD95 of 1.27 mm. These values closely approach the intra-rater variability of the more consistent human annotator (noting that this comparison involves different image samples). Using a very conservative threshold of 2 mm for the mean surface distance to define outliers, the in-distribution outlier rate was just 2%.
@@ Discussion
This work introduces a fully automated segmentation tool for brain dissection photographs, offering a practical and scalable alternative to resource-intensive annotation workflows. By leveraging the nnU-Net framework, our method achieves a level of accuracy that is comparable to that of expert human annotators. It yields strong performance on in-distribution data, and generalizes effectively to out-of-distribution images from a previously unseen site.
@@
To this end, future work may explore additional strategies to enhance robustness across domains. For example, synthetic data generation with domain randomization could further expose the model to a broader range of appearances and imaging conditions. Incorporating style transfer or domain adaptation techniques may also reduce sensitivity to site-specific artifacts. Another potential direction is to use lightweight image-level quality control (QC) mechanisms, such as binary pass/fail ratings, to complement traditional spatial metrics such as Dice and surface distance. Such QC frameworks may provide more intuitive assessments of segmentation reliability in large-scale, real-world workflows.
Ultimately, the practical impact of this tool lies in its potential to dramatically reduce the time burden of manual annotation. As we estimated, segmenting a typical dissection photograph takes approximately 10 minutes by hand. Extrapolated to our training dataset alone, this represents over 200 hours of expert time. Our tool can almost entirely remove this workload for most images and still substantially accelerate expert review when manual refinement is needed. This represents a significant step toward scalable, high-throughput analysis of neuropathology specimens, unlocking the potential of vast image archives that would otherwise be prohibitively time-consuming to label manually.
@@ Acknowledgments
This research was primarily funded by the National Institute of Aging (1R01AG070988). Additional support was provided by the National Institute of Mental Health (1RF1MH123195, 1UM1MH130981), the National Institute of Aging (1RF1AG080371), the National Institute of Biomedical Imaging and Bioengineering (1R01EB031114), and the National Institute of Neurological Disorders and Stroke (1R21NS138995). The UW BioRepository and Integrated Neuropathology (BRaIN) laboratory is supported by the National Institutes of Health (NIH) through the UW Alzheimer’s Disease Research Center (P30 AG066509), the Kaiser Permanente Washington Adult Changes in Thought (ACT) study (U19 AG066567), the BRAIN Initiative Cell Atlas Network (UM1MH134812), the Seattle Alzheimer’s Disease Brain Cell Atlas (U19AG060909), multiple collaborative projects supporting imaging, digital pathology, neuropathology, and related research tools (U24AG072458; U24NS133949; U24NS133945; U24NS135651; U01NS137500; and U01NS137484), the US Department of Defense (DoD W81XWH-21-S-TBIPH2), the Allen Institute for Brain Science, and the Nancy and Buster Alvord Endowment (to C.D.K.). Support for this research was provided in part by the BRAIN Initiative Cell Atlas Network (BICAN) grants U01MH117023, UM1MH134812 and UM1MH130981, the Brain Initiative Brain Connects consortium (U01NS132181, 1UM1NS132358-01), the National Institute for Biomedical Imaging and Bioengineering (1R01EB023281, R21EB018907, R01EB019956, P41EB030006), the National Institute on Aging (R21AG082082, 1R01AG064027, R01AG016495, 1R01AG070988), the National Institute of Mental Health (UM1MH130981, R01 MH123195, R01 MH121885, 1RF1MH123195), the National Institute for Neurological Disorders and Stroke , (1U24NS135561-01, R01NS070963, 2R01NS083534, R01NS105820, R25NS125599), and was made possible by the resources provided by Shared Instrumentation Grants 1S10RR023401, 1S10RR019307, and 1S10RR023043. Additional support was provided by the NIH Blueprint for Neuroscience Research (5U01-MH093765), part of the multi institutional Human Connectome Project. Much of the computation resources required for this research was performed on computational hardware generously provided by the Massachusetts Life Sciences Center (https:/ www.masslifesciences.com). The authors would like to thank the research participants and their families without whom this work would be impossible.
@@
Dr Hyman owns stock in Novartis; he serves on the SAB of Dewpoint and has an option for stock. He serves on a scientific advisory board or is a consultant for AbbVie, Alexion, Ambagon, Aprinoia Therapeutics, Arbor Bio, Arvinas, Avrobio, AstraZenica, Biogen, Bioinsights, BMS, Cell Signaling, Cure Alz Fund, CurieBio, Dewpoint, Etiome, Latus, Merck, Novartis, Paragon, Pfizer, Sanofi, Sofinnova, SV Health, Takeda, TD Cowen, Vigil, Violet, Voyager, WaveBreak. His laboratory is supported by research grants from the National Institutes of Health, Cure Alzheimer’s Fund, Tau Consortium, and the JPB Foundation – and sponsored research agreement from Abbvie and Sanofi. He has a collaborative project with Biogen and Neurimmune. Dr Fischl is an advisor to DeepHealth, a company whose medical pursuits focus on medical imaging and measurement technologies. BF’s interests were reviewed and are managed by Massachusetts General Hospital and Mass. General Brigham in accordance with their conflict of interest policies.